In research, it has been found that millions of bottled mineral water contain benzene (C6H6) at morality levels higher than concentrations considered lethal for rats. A sample of 250 mL mineral water has a concentration of 0.000151 M. What is the mass of benzene in the sample? Show your solutions

1
Expert's answer
2021-07-07T05:02:23-0400

number of moles =concentration"\times" volume"=0.000151\times 250/1000=3.775\times10^{-5}moles"

"C=12, H=1"

"RMM=12\times6+6\times 1=78"

"mass=RMM\times number\space of\space moles=78\times3.775\times10^{-5}=2.9445\times 10^{-3}g"